OpenStudy (watermelon14):

Consider the compound Al(NO3)3 and answer the following questions. What is the name of this compound? What is the molar mass of this compound? Explain how you calculated this value.

OpenStudy (watermelon14):

@Abhisar

OpenStudy (abhisar):

This compound is called as Aluminium Nitrate

OpenStudy (watermelon14):

yep i got that but how do you find the molar mass?

OpenStudy (abhisar):

Molar mass = Sum of molar masses of each component

OpenStudy (abhisar):

Got it ?

OpenStudy (watermelon14):

yep. Thank you:)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar In Al(NO\(_3\))\(_3\), mass of Al = 27, Molar mass of O=16, Molar mass of N=14. Molar mass of Al(NO\(_3\))\(_3\) = 27+(3\(\times\)14)+(9\(\times\)16)=212.99\(\approx\)213 \(\color{blue}{\text{End of Quote}}\)